# CloudCast – Total Cloud Cover Nowcasting with Machine Learning

### v1

@@ 3 Results
In this section we describe the results of the ablation study (Section 3-a), that were used to select the final model for further retraining. These results are discussed in depth in Section 3-b, and three case studies are presented in Section 3-c. Finally, the operational application is described briefly in section 3-d.
3.1 Ablation study
@@
To evaluate and score the models three different verification scores were used: mean absolute error (MAE), chi-squared value and power spectral density (PSD). The final score was the mean of the individual scores over grid and all the lead times. Chi-squared test compares the observed and expected frequencies across the dataset. The data was categorized into 20 bins and the chi-squared statistics for each model were compared to determine which models’ forecasts distribution aligns most closely with the ground truth. PSD was used to quantify the energy contained at different spatial scales, helping eliminate models that fail to capture small-scale features. Scale-separation was done using discrete fourier transform where the power of each frequency (scale) was the squared magnitude of the coefficients. Hanning windowing was used to prevent spectral leakage. Energy across different scales was averaged, placing greater emphasis on smaller scales since they inherently contain less energy.
Table 2: Results from the ablation study, sorted from best to worst. Each verification score was derived from the test dataset covering four months. Raw values are scaled between 0 and 1 using min-max normalization, where smaller value indicates better performance. The total score is the average of the three different tests. The best model for each category is highlighted in bold.
@@
mae-4-topo 0.000 0.390 0.0075 0.1310
mae-6-nd 0.037 0.379 0.0025 0.1380
mae-4-sun 0.000 0.418 0.0044 0.1395
mae-4-nd 0.074 0.355 0.0003 0.1417
mae-4-terrain 0.074 0.368 0.0128 0.1500
mae-2-nd 0.185 0.402 0.0001 0.1937
bcl1-4-nd 0.722 0.169 0.4646 0.4475
ssim-4-nd 0.426 1.000 0.0706 0.4939
mae-1-nd 1.000 0.521 0.0111 0.5055
bc-4-nd 0.907 0.044 0.7670 0.5671
logcosh-4-nd 0.796 0.182 0.7816 0.5806
mse-4-nd 0.926 0.000 1.0000 0.6356
@@
Results of the ablation study are presented in Table 2, ranking the models from best to worst. For brevity, only the average score across all seasons is shown. Models trained with MAE as loss function performed better than those with other types of loss functions. This can be partly attributed to the implicit advantage these models had, as they were trained and evaluated using the same metric. Nevertheless, these models also performed well on the other two tests. The model using MSE as loss function had most detail in the small scale structures, as indicated by the smallest value in PSD test, but at the same time it was the worst in chi-squared test, failing to reproduce the distribution of the ground truth. Based on the combined scores from the three tests, five out of the six best-performing models used MAE as the loss function and utilized four input images. The overall best-performing model was mae-4-dt, which besides the four input images also uses date and time information as predictors. This model was selected for further retraining.
Figure 4: Five-hour prediction from each of the models and the corresponding ground truth. Models are sorted from best to worst based on overall verification score. Forecast initialisation time is June 17 2024, 12 UTC.
Figure 4 presents a five-hour forecast from each model alongside the ground truth. Clear skies are shown in black, while overcast conditions are depicted in white. As expected, all forecasts appear blurrier compared to the ground truth. The models using binary cross-entropy, log-cosh, and MSE exhibit more broken cloud cover (grey areas) than the others. The model trained with the SSIM loss function stands out with high contrast, displaying good 0/1 separation, which is ideal; however, the quantitative verification scores did not support using SSIM as a loss function. A visual artifact in the top-right corner is due to limitations in the satellite instrument’s viewing angle.
3.2 Verification results
Based on the results of the ablation study MAE was chose as the loss function of CloudCast. Number of input images was four, and date and time were used as input features.
A useful forecast is one that has skill. To remove the possibility that a forecast is good by chance, the definition of a skilful forecast was adopted from Wilks (2011): the forecast outperforms a reference forecast, such as climatology or random forecast. Hogan et al. (2009) introduced mean absolute error skill score (MAESS) for continuous cloud cover data. MAESS compares a forecast against a reference forecast, typically either a random forecast or the climatological value. In this case, climatology was used: for each of the four verification months, a mean value was calculated for each grid point and used as the reference forecast. MAESS values range from zero or less (no skill, equal to or worse than climatology) to one (perfect forecast).
@@ 3.2 Verification results
MAESS was produced for each season individually and over all seasons. Model data was downscaled (MEPS) or upscaled (CloudCast) to match the resolution of the ground truth. In Figure 5 panel a), the MAESS values are plotted as a function of lead time over all four seasons. The persistence model achieves a perfect score for the 0-hour prediction, since it represents the last observed state before forecasting begins. However, its accuracy drops to half within the first hour and continues to decline as the forecast lead time increases. The MEPS model shows a consistently low MAESS value of around 0.3, indicating that the error characteristics are independent of the lead time. NWP models merge observational inputs, which are often spatially sparse and temporally irregular, with the background field (i.e., an earlier forecast) to create an internal grid-based representation of the atmospheric state. The significant difference in error characteristics between CloudCast and MEPS may suggest a notable discrepancy between cloudiness in the model space and in the observation space. CloudCast achieves a MAESS value of 0.83 at 0-hour, reflecting the penalty from its reduced horizontal resolution. Although forecast quality decreases with longer lead times, CloudCast’s decline is less steep than that of the persistence model. Overall, CloudCast outperforms all other models for all lead times except at 0-hour. Panel b) presents the MAESS scores for CloudCast across each distinct season. Spring consistently has the highest scores over all lead times, with a difference of about 0.1 compared to other seasons. Spring is characterised by dynamic small-scale processes initiated by increased solar radiation, leading to more variable cloud cover. Winter and summer exhibit very similar performance, which is somewhat surprising given their contrasting characteristics: winter is dominated by frequent low-pressure systems and limited short wave radiation, while summer, with its abundant daylight, promotes convective activity, leading to daytime cloud development and clearer conditions at night. The lowest scores are observed in autumn, which, like winter, tends to have more overcast conditions. Panel c) displays the MAESS for intra-hour predictions over all seasons, using EXIM as the benchmark model. At 0-hour CloudCast displays the same bias observed in panel a), attributed to its coarser horizontal resolution. EXIM’s performance declines noticeably by the 15-minute mark, where it scores lower than the persistence model. This downward trend continues, albeit at a slower rate, throughout the forecast period. CloudCast consistently outperforms both EXIM and the persistence model across all forecast lead times, with a margin of approximately 0.1.
@@
Figure 6 shows the mean bias of CloudCast, defined as the difference between forecast and observation, as a function of lead time. CloudCast tends to overestimate cloud cover, reflected by a positive bias. The bias increases linearly across all seasons, becoming more pronounced with longer lead times, reaching a mean bias of 0.03 by forecast hour 5. Winter and summer generally exhibit less bias compared to spring and autumn.
In figure 7 the seasonal spatial distribution of the bias is displayed on a map centered around Finland and its neighbouring areas. Positive bias is the dominant pattern across all four seasons, especially in autumn and winter. Negative bias becomes more prevalent in spring and summer, with the summer showing the most pronounced balance between positive and negative biases. The northern and western regions consistently show more overestimation, while the eastern and southern parts exhibit more underestimation, particularly in spring and summer. Topography or land-sea distribution appears to have no significant influence on the bias.
@@
The Fractions Skill Score (FSS) (Roberts and Lean, 2008) is a widely adopted metric for spatial verification of categorized events. As a neighbourhood-based method, FSS evaluates how well a forecasted event matches the observed event within a specified spatial threshold. It is commonly used to quantify the spatial scale at which the forecast remains useful or skilful (Mittermaier and Roberts, 2010).
@@
Figure 8: Fractions Skill Score over all four seasons. Limits for skilful (uniform) and random (random) forecasts are drawn. Higher values indicate better performance.
The four panels in Figure 8 display the Fractions Skill Score (FSS) for the clear, partly cloudy, mostly cloudy, and overcast categories. Scores are calculated for all seasons, with forecasts made in 15-minute increments. The threshold for a skillful forecast, referred to as “uniform” by Roberts and Lean (2008), is represented by a dashed line. Panels b) and c) also include a second line indicating the threshold for a random forecast. In the clear sky category, CloudCast remains skillful for up to three hours, even at the smallest spatial scales. As the scale increases to 30 km, the forecast remains skillful for up to five hours. At the widest scale of 1970 km, the FSS reaches a perfect score of 1 for up to three hours. Due to computational constraints, the FSS was not calculated for the entire domain (2300 km), but based on current data, it is likely the FSS would fall short of 1 at five hours, suggesting increased bias at longer lead times. Panels b) and c) show the FSS for the partly cloudy and mostly cloudy categories, where the forecasts are less accurate compared to clear skies. For partly cloudy conditions, the forecast is skillful up to three hours at a spatial scale of about 70 km, and for mostly cloudy, up to 40 km. However, after 30 minutes for partly cloudy and 90 minutes for mostly cloudy, predictions fall below the random forecast threshold at the smallest scales. In the overcast category, the skillful range starts at around 150 minutes for the 5 km scale, reaching three hours at the 10 km scale. The FSS reaches a perfect value for the longest forecast at scales around 400 km, indicating less bias for this category compared to clear skies.
@@
Figure 9 displays the row-normalized confusion matrices for different cloudiness categories, comparing CloudCast (left) and MEPS (right). For CloudCast, the diagonal is well-represented, indicating a strong alignment between forecasted and observed categories. Specifically, 77% of clear sky conditions and 61% of overcast conditions were accurately predicted. However, there is a slight tendency to overpredict mostly cloudy as overcast (38%). The largest discrepancy appears in the partly cloudy category, where only 31% of forecasts were accurate, partly due to the limited data available for this category. Multi-category errors, represented in the corners of the matrix, are minimal. MEPS, on the other hand, tends to overpredict mostly cloudy (74%) and partly cloudy (49%) conditions. Overall, CloudCast provides a more balanced prediction across categories, while MEPS struggles particularly with clear and partly cloudy conditions.
@@ 3.3 Case studies
Figure 10: Five hour cloud cover forecast initialized at July 10 2024 12 UTC. Top row is the observed value, middle row CloudCast forecast and bottom row MEPS forecast.
@@
Figure 11: Five hour cloud cover forecast initialized at August 9 2024 06 UTC. Top row is the observed value, middle row CloudCast forecast and bottom row MEPS forecast.
@@
The third case study is presented in Figure 12. The intra-hour forecasts from CloudCast and EXIM are shown in the middle and bottom rows, respectively. EXIM does not include the initialization time data. Over the span of 60 minutes, little change is observed, especially given the absence of small-scale dynamics in the ground truth. CloudCast captures the large-scale structure of the cloud cover but produces forecasts that appear blurry. As in the first case study, CloudCast tends to underpredict areas of overcast cloud. EXIM, on the other hand, retains small-scale features but lags in cloud development and fails to capture the full extent of cloud cover, particularly in the longer lead times.
Figure 12: Intra-hour cloud cover forecast initialized on August 30 2024 11 UTC. Top row is the observed value, middle row CloudCast forecast and bottom row EXIM forecast.
@@ 3.4 Operational application
CloudCast has been used to produce the total cloud cover forecast to FMI’s operational nowcasting system (Hieta et al., 2021) since January 2023. A five hour forecast is produced every fifteen minutes, and the last two hours of the forecast are used to blend it with NWP model forecast. The input effective cloudiness data is preprocessed to reduce the defects inherent in the data itself. The operational forecasting process, including reading input data from object storage, evaluating the model, upscaling the results to 2.5km resolution, converting the output data into GRIB file format, and writing the output data takes around 10 seconds using an Nvidia MIG 24G GPU.
4 Discussion

### v2

@@ 3 Results
In this section we describe the results of the sensitivity experiments (Section 3a), that were used to select the final model for further retraining. These results of the final model are discussed in Section 3b, and three case studies are presented in Section 3c. Finally, the operational application is described briefly in section 3d.
3.1 Sensitivity experiments
@@
To evaluate and score the models three different verification scores were used: mean absolute error (MAE), chi-squared value and power spectral density (PSD). The final score was the mean of the individual scores over grid and all the lead times. Chi-squared test compares the observed and expected frequencies across the dataset. The data was categorized into 20 bins and the chi-squared statistics for each model were compared to determine which models’ forecasts distribution aligns most closely with the ground truth. PSD was used to quantify the energy contained at different spatial scales, helping eliminate models that fail to capture small-scale features. Scale-separation was done using discrete fourier transform where the power of each frequency (scale) was the squared magnitude of the coefficients. Hanning windowing was used to prevent spectral leakage that otherwise would distort the results for the smallest and largest scales. Power density values were averaged across scales, with each scale weighted by the inverse of its length to emphasize the smaller scales more strongly.
Table 2: Results from the sensitivity experiments, sorted from best to worst. Each verification score was derived from the test dataset covering four months. Models are ranked from best to worst (best=1, worst=15) ie. smaller value indicates better performance. The total rank is derived by summing of the points from the three different tests and sorting from smallest to largest. The best model for each category is underlined.
@@
mae-6-dt 1 3 5 2
mae-2-dt 6 4 1 3
mae-4-sun 3 5 7 4
mae-6-nd 5 6 6 5
mae-4-topo 3 7 8 6
mae-2-nd 9 8 3 7
mae-4-nd 6 10 4 7
mae-4-terrain 6 9 10 9
ssim-4-nd 15 1 11 10
bcl1-4-nd 10 12 12 11
mae-1-nd 14 11 9 11
bc-4-nd 11 14 13 13
logcosh-4-nd 12 13 14 14
mse-4-nd 13 15 15 15
@@
Results of the sensitivity experiments are presented in Table 2, ranking the models from best to worst. For brevity, only the total score over all seasons is shown. Models trained with MAE as loss function performed better than those with other types of loss functions. This can be partly attributed to the implicit advantage these models had, as they were trained and evaluated using the same metric. Nevertheless, these models also performed well on the other two tests. The model using MSE as loss function had most detail in the small scale structures, as indicated by the smallest value in PSD test, but at the same time it was the worst in chi-squared test, failing to reproduce the distribution of the ground truth. Based on the combined scores from the three tests, five out of the six best-performing models used MAE as the loss function and used four input images. The overall best-performing model was mae-4-dt, which besides the four input images also uses date and time information as predictors. This model was selected for further retraining in horizontal resolution of 5km.
Figure 4: Five-hour prediction from each of the models and the corresponding ground truth. Models are sorted from best to worst based on overall verification score. Forecast initialisation time is June 17 2024, 1200 UTC.
Figure 4 presents a five-hour forecast from each model alongside the ground truth. Clear skies are shown in white, while overcast conditions are depicted in black. As expected, all forecasts appear blurrier compared to the ground truth. The models using binary cross-entropy, log-cosh, and MSE exhibit more broken cloud cover (grey areas) than the others. The model trained with the SSIM loss function stands out with high contrast, displaying good 0/1 separation, which is ideal; however, the quantitative verification scores did not support using SSIM as a loss function. A visual artifact in the top-right corner is due to limitations in the satellite instrument’s viewing angle.
3.2 Full resolution model
Based on the results of the sensitivity experiments MAE was chosen as the loss function of CloudCast. The number of input images was four, and date and time were used as input features.
A useful forecast is one that has skill. To remove the possibility that a forecast is good by chance, the definition of a skillful forecast was adopted from Wilks (2011): the forecast outperforms a reference forecast, such as climatology or random forecast. Hogan et al. (2009) introduced mean absolute error skill score (MAESS) for continuous cloud cover data. MAESS compares a forecast against a reference forecast, typically either a random forecast or the climatological value. In this case, climatology was used: for each of the four verification months, the mean value was calculated for each grid point and used as the reference forecast. MAESS values range from zero or less (no skill, equal to or worse than climatology) to one (perfect forecast).
@@ 3.2 Full resolution model
The models and ground truth each had different horizontal resolutions. We performed nearest point interpolation to match the models to the ground truth. For MEPS, we upscaled the data from 2.5km to 3km, for cloudcast we downscaled from 5km to 3km. EXIM data was already produced in the same resolution as ground truth. MAESS was computed separately for each season and across all seasons. Figure 5a shows MAESS as a function of lead time, averaged over all seasons (spring: MAM, summer: JJA, autumn: SON, winter: DJF). The persistence model achieves a perfect score for the 0-hour prediction, since it represents the last observed state before forecasting begins. However, its accuracy drops to half within the first hour and continues to decline as the forecast lead time increases. The MEPS model shows a consistently low MAESS value of around 0.3, indicating that the forecast skill is independent of the lead time. NWP models merge observational inputs, which are often spatially sparse and temporally irregular, with the background field (i.e., an earlier forecast) to create an internal grid-based representation of the atmospheric state. The significant difference in forecast skill between CloudCast and MEPS may suggest a notable discrepancy between cloudiness in the model space and the observation space. CloudCast achieves a MAESS value of 0.83 at 0-hour, a limitation due to its lower horizontal resolution. Although forecast quality decreases with longer lead times, CloudCast’s decline is less steep than that of the persistence model. Overall, CloudCast outperforms all other models for all lead times except at 0-hour. Panel b) presents the MAESS scores for CloudCast across each distinct season. Spring consistently has the highest scores over all lead times, with a difference of about 0.1 compared to other seasons. Spring is characterised by dynamic small-scale processes initiated by increased solar radiation, leading to more variable cloud cover. Winter and summer exhibit very similar performance, which is somewhat surprising given their contrasting characteristics: winter is dominated by frequent low-pressure systems and limited short wave radiation, while summer, with its abundant daylight, promotes convective activity, leading to daytime cloud development and clearer conditions at night. The lowest scores are observed in autumn, which, like winter, tends to have more overcast conditions. Panel c) displays the MAESS for intra-hour predictions over all seasons, using EXIM as the benchmark model. At 0-hour CloudCast displays the same bias observed in panel a), attributed to its coarser horizontal resolution. EXIM’s performance declines noticeably by the 15-minute mark, where it scores lower than the persistence model. This downward trend continues, albeit at a slower rate, throughout the forecast period. CloudCast consistently outperforms both EXIM and the persistence model across all forecast lead times, with a margin of approximately 0.1. At all lead times, EXIM underperforms compared to the persistence forecast — likely because it tends to shift clouds in the wrong direction. Given that cloud movement is generally slow, even a static persistence forecast remains a strong baseline.
@@
Figure 6 shows the mean bias of CloudCast, defined as the difference between forecast and observation, as a function of lead time. CloudCast tends to overestimate cloud cover, reflected by a positive bias. The bias increases linearly across all seasons, becoming more pronounced with longer lead times, reaching a mean bias of 0.03 by forecast hour 5. Winter and summer generally exhibit less bias compared to spring and autumn. The positive bias may stem from the model’s difficulty in capturing cloud dissipation processes over time. Additionally, as forecast uncertainty increases, the model expresses this as a blurring effect, leading to higher bias. This occurs because cloud cover data is typically concentrated near zero and one, meaning that a gradual shift toward intermediate values inflates the mean bias.
In figure 7 the seasonal spatial distribution of the bias is displayed on a map centered around Finland and its neighbouring areas. Positive bias is the dominant pattern across all four seasons, especially in autumn and winter. Negative bias becomes more prevalent in spring and summer, with the summer showing the most pronounced balance between positive and negative biases. The northern and western regions consistently show more overestimation, while the eastern and southern parts exhibit more underestimation, particularly in spring and summer. Topography or land-sea distribution appears to have no significant influence on the bias. During winter low sun angle and snow covered ground can alter surface reflectance and emissivity, potentially affecting the data used for prediction. In summer, high radiance levels may lead to nonlinear cloud formation and dissipation processes that the model struggles to capture accurately. While the diurnal cycle is not explicitly represented in the data, they may still contribute to seasonal biases.
@@
The Fractions Skill Score (FSS) (Roberts and Lean, 2008) is a widely adopted metric for spatial verification of categorized events. As a neighbourhood-based method, FSS evaluates how well a forecasted event matches the observed event within a specified spatial threshold. It is commonly used to quantify the spatial scale at which the forecast remains useful or skillful (Mittermaier and Roberts, 2010).
@@
Figure 8: Fractions Skill Score over all four cloudiness categories. Limits for skillful and random forecasts are drawn. A skillful forecast is better than a reference forecast; in the case of FSS a reference forecast is a forecast that assigns the overall observed frequency of the event uniformly across the entire domain. In the case of the random forecast the frequencies are assigned based on the climatology.
The four panels in Figure 8 display the Fractions Skill Score (FSS) for the clear, partly cloudy, mostly cloudy, and overcast categories. Scores are calculated for all seasons, with forecasts made in 15-minute increments. The threshold for a skillful forecast, referred to as “uniform” by Roberts and Lean (2008), is represented by a dashed line. Panels b) and c) also include a second line indicating the threshold for a random forecast. In the clear sky category, CloudCast remains skillful for up to three hours, even at the smallest spatial scales. As the scale increases to 30 km, the forecast remains skillful for up to five hours. At the widest scale of 1970 km, the FSS reaches a perfect score of 1 for up to three hours. Due to computational constraints, the FSS was not calculated for the entire domain (2300 km), but based on current data, it is likely the FSS would fall short of 1 at five hours, suggesting increased bias at longer lead times. Panels b) and c) show the FSS for the partly cloudy and mostly cloudy categories, where the forecasts are less accurate compared to clear skies. For partly cloudy conditions, the forecast is skillful up to three hours at a spatial scale of about 70 km, and for mostly cloudy, up to 40 km. However, after 30 minutes for partly cloudy and 90 minutes for mostly cloudy, predictions fall below the random forecast threshold at the smallest scales. In the overcast category, the skillful range starts at around 150 minutes for the 5 km scale, reaching three hours at the 10 km scale. The FSS reaches a perfect value for the longest forecast at scales around 400 km, indicating less bias for this category compared to clear skies. Overall the FSS results indicate that CloudCast provides the most skillful forecasts for clear skies and overcast conditions, while performance degrades more quickly for partly and mostly cloudy cases. The model retains skill for up to three hours in fine-scale forecasts and up to five hours at larger scales, with the strongest performance in stable cloud regimes. However, skill drops rapidly for mixed cloud conditions, aligning with the increasing bias seen in Figure 6. This suggests that uncertainty in cloud evolution limits forecast accuracy.
@@
Figure 9 displays the row-normalized confusion matrices for different cloudiness categories, comparing CloudCast (left) and MEPS (right). For CloudCast, the diagonal is well-represented, indicating a strong alignment between forecasted and observed categories. Specifically, 77% of clear sky conditions and 61% of overcast conditions were accurately predicted. However, there is a slight tendency to overpredict mostly cloudy as overcast (38%). The largest discrepancy appears in the partly cloudy category, where only 31% of forecasts were accurate, partly due to the limited data available for this category. Multi-category errors, represented in the corners of the matrix, are minimal. MEPS, on the other hand, struggles in clear, partly cloudy, and mostly cloudy conditions. Overall, CloudCast provides a more balanced prediction across categories, while MEPS struggles particularly with clear and partly cloudy conditions.
@@ 3.3 Case studies
Figure 10: Five hour cloud cover forecast initialized at July 10 2024 1200 UTC. A typical summertime scenario is illustrated, where Finland is covered by convection-initiated, small-scale cloud patterns (large cumulus or stratocumulus). As the day progresses, the small-scale clouds gradually dissipate, while a frontal-based system advances towards northeast, covering the western part of Finland by the end of the forecast period. Top row is the observed value, middle row CloudCast forecast and bottom row MEPS forecast.
@@
Figure 11: Five hour cloud cover forecast initialized at August 9 2024 0600 UTC. The western part of the domain is initially covered by clouds, while the eastern region experiences clearer skies. Gradually clouds begin to cluster over eastern Finland, eventually spreading to cover the entire country. Top row is the observed value, middle row CloudCast forecast and bottom row MEPS forecast.
@@
The third case study is presented in Figure 12. The intra-hour forecasts from CloudCast and EXIM are shown in the middle and bottom rows, respectively. Over the span of 60 minutes, little change is observed, largely due to the lack of small-scale cloud evolution in the ground truth. A frontal cloud covers the western half of the domain, while Southern Finland remains under clear skies. Cloud development is minimal, with a single larger cloud element over Eastern Finland gradually dissipating. CloudCast captures the large-scale structure of the cloud cover but produces forecasts that appear blurry. As in the first case study, CloudCast tends to underpredict areas of overcast cloud. EXIM, on the other hand, retains small-scale features but lags in cloud development and fails to capture the full extent of cloud cover, particularly in the longer lead times.
Figure 12: Intra-hour cloud cover forecast initialized on August 30 2024 1100 UTC. A frontal cloud covers the western half of the domain, while Southern Finland remains under clear skies. Cloud development is minimal, with a single larger cloud element over Eastern Finland gradually dissipating. Top row is the observed value, middle row CloudCast forecast and bottom row EXIM forecast.
@@ 3.4 Operational application
CloudCast has been used to produce the total cloud cover forecast to FMI’s operational nowcasting system (Hieta et al., 2021) since January 2023. A five hour forecast is produced every fifteen minutes, and the last two hours of the forecast are used to blend it with NWP model forecast. The input effective cloudiness data is preprocessed to reduce the defects inherent in the data itself. However, since the operational forecast began after the model training was completed, we have not directly quantified the improvement in forecast skill relative to the verification results in Figure 5. Additionally, verification is challenging because the ground truth data is also preprocessed, meaning that the improved quality is not necessarily reflected in simple objective scores like MAE but requires either verification against third-party observations or subjective evaluation by domain experts.
The operational forecasting process, including reading input data from object storage, evaluating the model, downscaling the results to 2.5km resolution, converting the output data into GRIB file format, and writing the output data takes around 10 seconds using an Nvidia MIG 24G GPU.
4 Summary and discussion